# SAP-CoPE: Social-Aware Planning using Cooperative Pose Estimation with Infrastructure Sensor Nodes

### v1

@@ IV-B Delay-aware Global Perception
The delay-aware global perception framework at the CL is designed to overcome the challenges posed by communication delays in highly dynamic indoor environments. Each sensor node assigns timestamps to detected objects, enabling the system to compensate for delays by comparing these timestamps to the current time at the central node. Using timestamp-based delay compensation, the system adjusts object positions through motion prediction models, ensuring real-time environmental awareness. For human tracking, a non-linear motion model is used to predict movement direction and speed changes. This predictive capability improves the system’s ability to maintain consistent and reliable tracking across different ISNs.
After compensating for delays, the system employs a weighted fusion strategy to combine the local perception results from multiple ISNs. This approach ensures a cohesive and accurate representation of the environment, even when data arrives asynchronously from different sources. By combining advanced local and global perception techniques, the system achieves real-time detection and tracking, effectively compensating for delays while optimizing multi-sensor data fusion in complex indoor environments.
V Integrated Motion Planning and Tracking
@@ V-A Kinematic-based Motion Model
$\dot{v_{f}}=a_{f},\quad\dot{v_{r}}=a_{r}$ (5)
$v_{c}=\frac{v_{f}\cos(\delta_{f})+v_{r}\cos(\delta_{r})}{2\cos(\beta)}$ (7)
Where the $X$, $Y$, and $\psi$ represent robot’s pose in the global frame; $\beta$ is the side-slip angle of C.G., $v_{c}$ is the C.G. velocity; $l_{f}$ and $l_{r}$ are the distances from the C.G. to the front and rear wheels; and $v_{f}$, $v_{r}$, $a_{f}$, $a_{r}$, $\delta_{f}$, and $\delta_{r}$ are the velocity, acceleration, and steering of the front and rear wheels, respectively. In this paper, the system states, $\xi$, system outputs, $\eta$, and system inputs, $u_{c}$ are defined as follows.
$\zeta=\eta=\begin{bmatrix}X&Y&\theta&v_{f}&v_{r}\end{bmatrix}^{T}$ (8)
$u_{c}=\begin{vmatrix}a_{f}&a_{r}&\delta_{f}&\delta_{r}\end{vmatrix}^{T}$ (9)
V-B Personal Space Model
@@
Public distance $>$ 3.5 m Public speaking
@@
This paper adapts the concept of PS[25] to model human behavior in interaction with autonomous systems. Personal space captures an individual’s non-verbal and non-contact communication channels, influencing how they perceive proximity to others. Based on the studies from Edward T. Hall [29] and R. Sommer [30], PS can be divided into four regions: public distance, social distance, personal distance, and intimate distance. These regions define the various zones of comfort and social interaction. The graphical representation and the relative distances can be found in Fig. 4 and Table I.
@@
Figure 4: Graphical Representation of Personal Field. a) Definition of Personal Field in psychology [25]. b) Constructed Personal Field using Mathematical Expression
@@
Let person $P$ is at position $\{x_{p},y_{p}\}$ in the global frame with heading angle $\theta_{p}$; and let another person $Q$ is at position $\{x_{q},y_{q}\}$ in the global frame. The PS model of person $P$ relative to person $Q$ is denoted as $\Omega$, which is divided into two parts, where $\Omega_{f}$ represents the front PS, and the $\Omega_{r}$ is the rear PS. Additionally, $\gamma(x_{p})=1$ if $x_{p}\geqslant 0$ and $\gamma(x_{p})=0$ otherwise. The expression of PS model can be found below.
$\Omega=\gamma(x_{p})\Omega_{f}+(1-\gamma(x_{p}))\Omega_{r}$ (10)
@@
Each front and rear personal space $\Omega_{i},i\in\{f,r\}$ is defined as a 2D Gaussian function using the covariance matrix $\Sigma$, relative distance $d$, and heading angle $\theta$.
$\Omega_{i}=e^{-0.5d^{T}\Sigma_{i}^{-1}d},i\in\{f,r\}$ (11)
$d=\begin{bmatrix}x_{p}-x_{q}\\ y_{p}-y_{q}\end{bmatrix}\begin{bmatrix}cos(\theta_{p})&sin(\theta_{p})\\ -sin(\theta_{p})&cos(\theta_{p})\end{bmatrix}$ (12)
@@
$\Sigma$ defines the shape of the PS. Based on the studies of Shozo[31], it is suggested that the front PS of a human should be twice as large as the rear, left, and right sides. To implement this, the $\Sigma_{f}$ and $\Sigma_{r}$ are defined as follows, where $\sigma_{xx}=2\sigma_{yy}$.
@@
Amaoka et al. [25] stated that the values of $\sigma_{xx}$ should depend on several factors such as age, gender, etc.; however, since this paper is focused on the overall framework for indoor mobility system, the values of $\sigma_{xx}$ is assumed to be fixed at $0.5$ for all people when conducting simulations and experiments.
V-C Model Predictive Controller Design
The formulation of the proposed integrated controller can be expressed using the following equations. Equ. 14 is the cost function of the MPC; Specifically, $J_{1}$ is the APF function for non-human-type obstacles, with the detailed expression provided in our previous work [7, 32]; $J_{2}$ is the PS model for humans; $J_{3}$ and $J_{4}$ are the tracking and control costs, respectively, which are the same as in regular MPC.
Furthermore, Equ. 18 and 19 describe the motion model of the robot; Equ. 20 defines the control rate constraint; Equ. 21 presents the control constraints; Equ. 22 specifies the state constraints; and lastly, Equ. 23 defines the wheel speed constraint to prevent wheel slippage caused by speed differences between the front and rear wheels. More details can be found in [7].
The Taylor series expansion approximates the PS and APF functions into their quadratic form, where the nominal points are the predicted position of the robot, human, and obstacles using the Constant Velocity and Turning Rate (CVTR) over $N_{p}$ time steps. The optimization problem is solved using OSQP[33].
$\arg\min_{u_{c}}J=J_{1}+J_{2}+J_{3}+J_{4}$ (14)
@@ V-C Model Predictive Controller Design
$J_{1}=\sum_{p=1}^{N_{obs}}APF(p),J_{2}=\sum_{q=1}^{N_{human}}PS(q)$ (15)
$J_{3}=\sum_{i=0}^{N_{p}-1}\|\eta(k+i+1)-\eta_{\text{Ref}}(k+i+1)\|_{Q}^{2}$ (16)
$J_{4}=\sum_{j=0}^{N_{c}-1}\|\Delta u_{c}(k+j)\|^{2}_{R}$ (17)
@@
$\zeta(k+i+1)=\tilde{A}_{k}\zeta(k+i)+\tilde{B}_{k}\Delta u_{c}(k+j)+\tilde{d}_{k}(k+i)$ (18)
$\eta(k+i+1)=\tilde{C}_{k}\zeta(k+i+1)$ (19)
@@
$\eta_{\min}\leq\eta(k+i)\leq\eta_{\max}$ (22)
$\left|E_{K}\Delta u_{c}(k+j)+g(\zeta_{0}(k),u_{0}(k))\right|\leq 0.1$ (23)
@@
VI Simulation
The parameters for the robot, perception system, and integrated controller are summarized in the table below. TThe proposed perception system and controller were implemented and tested in both simulation and experiments using the Python platform.
TABLE II: Robot and MPC Tuning Parameters
Symbol Value Symbol Value
$l_{f},l_{r}$ 1.2 m $N_{p}$ 20
$u_{c\max}$ $\left[1,1,\frac{\pi}{2},\frac{\pi}{2}\right]$ $N_{c}$ 10
$u_{c\min}$$\left[1,1,\frac{\pi}{2},\frac{\pi}{2}\right]$ $Q$ $diag(2,2,6,10,10)$
$\Delta u_{c\max}$ $\left[0.8,0.8,\frac{\pi}{12},\frac{\pi}{12}\right]$ $R$ $100~{}diag(3,3,4,4)$
$\Delta u_{c\min}$$\left[0.8,0.8,\frac{\pi}{12},\frac{\pi}{12}\right]$ $T_{s}$ 0.1 sec
$\eta_{\max}$ $\left[\text{Inf},\text{Inf},\text{Inf},1.0,1.0\right]$
$\eta_{\min}$$\left[\text{Inf},\text{Inf},\text{Inf},0.1,0.1\right]$
Figure 5: Simulation Results Comparison between Baseline and Proposed Method
Fig. 5 demonstrates the results of the baseline algorithm, which models all obstacles using the APF, and the proposed SAP-CoPE approach. In this simulation, the robot encounters two obstacles: a static obstacle represented as a green rectangle and a human represented as a green circle with the black line indicating their heading angle. The robot’s task is to follow the reference path, defined by the gray dashed line, while avoiding the obstacles.
In the figure, the baseline algorithm successfully avoids the collision with both obstacles while staying close to the reference path, which can be considered a successful outcome; however, when encountering the human-type obstacle, the robot enters the human’s PS after making the 90-degree turn. This may disturb the human and introduce psychological pressure, as well as increase the risk of collision.
In contrast, the SAP-CoPE, which incorporates the personal space model, can avoid collisions and maintain an appropriate distance from humans. It is worth noting that the robot follows the same trajectory when encountering a stationary obstacle as the baseline algorithm, with deviations occurring only when encountering human-type obstacles. This demonstrates the effectiveness of the presented PS model in maintaining human comfort.
VII Experiments
VII-A Perception Experiments
VII-A1 Evaluation Metrics
@@
VII-A2 Results
We compare two pose estimation methodologies: Camera Pose Estimation uses only image-based detections, and Camera-LiDAR Pose Estimation incorporates matched LiDAR points for each joint, if available. Quantitative results are summarized in Table III, while spatial distributions are provided in Figs. 6-7.
TABLE III: Pose Estimation Performance Evaluation

### v2

@@ IV-B Delay-aware Global Perception
In ISNs system, computational and communication delays can introduce inconsistencies in perception, severely degrading the performance of following motion planning task. To address this, our previous work proposed a delay-aware cooperative perception pipeline that effectively mitigates latency effects and provides consistent, real-time perception outputs [23]. It contains three key components: global fusion across sensor nodes, multi-object tracking and delay compensation.
Each ISN sends its local perception results to the CL, and the detections are matched based on spatial proximity and class labels. For the objects detected by multiple ISNs, instead of simply averaging their positions like previous works [23], it refines the 3D human pose estimation via Eqn. 6 at the CL by fusing multi-view pose constraints from different ISNs to enhance robustness against occlusions and partial views.
Then, it performs multi-object tracking using an extended Kalman filter (EKF) for each detected object to keep track of their states over time. Human trajectories, in particular, are modeled using a non-linear dynamic system capable of capturing changes in walking direction and speed, which enhances robustness in crowded and dynamic environments.
To further improve real-time responsiveness, the system incorporates timestamp-based delay compensation. Each ISN assigns a local timestamp to detected objects, allowing the CL module to compute the delay relative to its current time. Object positions are then extrapolated forward using motion prediction models, effectively aligning stale data with the current system state. This compensation mechanism ensures that the fused perception remains temporally consistent, even under variable network or processing latencies.
V Socially-aware Motion Planning
@@ V-A Kinematic-based Motion Model
Variables $X$ and $Y$ represent robot’s position in the ground-fixed frame; $\psi$ is the robot’s yaw angle; $\beta$ is the side-slip angle of C.G.; $v_{c}$ is the C.G. velocity; $l_{f}$ and $l_{r}$ are the distances from the C.G. to the front and rear wheels; and $v_{f}$, $v_{r}$, $\delta_{f}$, and $\delta_{r}$ are the velocity and steering of the front and rear wheels, respectively. In this paper, the system states, $\xi$, system outputs, $\eta$, and system control inputs, $u_{c}$ are defined as follows.
$\zeta=\eta=\begin{bmatrix}X&Y&\psi\end{bmatrix}^{T}$ (11)
$u_{c}=\begin{bmatrix}v_{f}&v_{r}&\delta_{f}&\delta_{r}\end{bmatrix}^{T}$ (12)
V-B Personal Space Field
@@
Public distance >> 3.5 m Public speaking
@@
This paper adapts the concept of personal space (PS)[2] to model human behaviour in interaction with the robot. Compare using the APF to represent the human, the PS model captures an individual’s non-verbal and non-contact communication channels, influencing how they perceive proximity to others. Based on the studies from Edward T. Hall [10] and R. Sommer [35], PS can be divided into four regions: public distance, social distance, personal distance, and intimate distance. These regions define the various zones of comfort and social interaction. The graphical representation and the relative distances can be found in Fig. 4 a) and Table I.
By assigning appropriate values to each region, as shown in Fig. 4 b), the robot is able to maintain a psychologically safe distance from nearby humans. In this work, we adopt the asymmetric shaped field proposed in [2], where the field is intentionally shaped to be nonuniform in the y-direction and concentrated along the human’s facing orientation. This design reflects the natural psychological comfort zones observed in human interactions, where individuals are more sensitive to being disturbed from the front. As a result, the robot can anticipate and steer earlier to avoid potential discomfort.
In contrast, the conventional SFM and Artificial Potential Field (APF) are normally symmetrical in both $x$ and $y$ directions, which requires enlarging the overall field size to achieve a similar early-dodging behavior from the robot. However, because the conventional field covers an equal area in front of and behind the human, it can lead to inefficient navigation behavior, such as unnecessary deceleration or even being trapped when operating in densely populated environments. The proposed PS field, therefore, enables more natural and efficient avoidance while ensuring safe interpersonal distances.
@@
Figure 4: Graphical Representation of Personal Space Field. a) Definition of Personal Field in psychology [2]. b) Constructed Personal Field using Mathematical Expression. In both figures, $X$ and $Y$ represent the body-fixed coordinate of the human.
@@
The formulation of the PS field can be found in the following. Let person $P$ with pose $\{x_{p},y_{p}\}$ at the position in the ground-fixed frame with heading angle $\theta_{p}$; and let another person $Q$ at position of $\{x_{q},y_{q}\}$ in the ground-fixed frame. The PS field of person $P$ relative to person $Q$ is denoted as $\Omega_{pq}$, which is divided into two parts, where $\Omega_{pq,f}$ represents the front PS, and $\Omega_{pq,r}$ is the rear PS.
$\Omega_{pq}=\gamma\Omega_{pq,f}+(1-\gamma)\Omega_{pq,r}$ (13)
Where, $\gamma$ is a function of $x_{p}$, such that $\gamma(x_{p})=1$ if $d\geqslant 0$ and $\gamma(x_{p})=0$ otherwise. This relationship can be approximated by using the tanhtanh function in Equ. 14, where $d$ is the relative position of person $Q$ to person $P$ in the body coordinate of person $P$, which can be expressed in Equ. 15. The variable $k$ is a constant that controls the slope of the tanhtanh’s curve. To allow the $\gamma$ becomes to 1 immediately when $d$ is 0, we assigned a large value to $k$.
$\gamma(x_{p})=0.5(tanh(k\cdot d/2)+1)$ (14)
$d=\begin{bmatrix}\cos\theta_{p}&\sin\theta_{p}\\ -\sin\theta_{p}&\cos\theta_{p}\end{bmatrix}\begin{bmatrix}x_{q}-x_{p}\\ y_{q}-y_{p}\end{bmatrix}$. (15)
@@
On the other hand, each front and rear personal space $\Omega_{pq,i},i\in\{f,r\}$ is defined as a 2D Gaussian function using the covariance matrix $\Sigma$ and relative distance $d$.
$\Omega_{pq,i}=e^{-0.5d^{T}\Sigma_{i}^{-1}d},i\in\{f,r\}$ (16)
@@
Where, $\Sigma_{i}$ defines the shape of the $\Omega_{pq,i}$. Based on the studies of Shozo[32], it is suggested that the front PS, $\Omega_{pq,f}$, of a human should be twice as large as the rear, left, and right sides. To implement this, the $\Sigma_{f}$ and $\Sigma_{r}$ are defined as follows, where $\sigma_{xx}=2\sigma_{yy}$.
@@
Amaoka et al. [2] stated that the values of $\sigma_{xx}$ should depend on several factors such as age, gender, etc.; however, since this paper is focused on the overall framework for indoor mobility system, the values of $\sigma_{xx}$ is assumed to be fixed for all people when conducting experiments.
V-C Model Predictive Control-based Planner Design
To achieve real-time computational efficiency, both the nonlinear kinematic motion equations and the PS field were approximated using the Taylor series. Specifically, the nominal points for the kinematic model were selected as the current state and control inputs measurements. For the PS field, the nominal points were defined as the predicted positions of the humans over the prediction horizon, $N_{P}$, obtained using the Constant Velocity and Turning Rate (CVTR) model. This approximation ensures the optimization problem is quadratic, allowing it to be efficiently solved using the OSQP solver [36].
The proposed MPC consists of three objective terms and a set of constraints. Equ. 18 is the cost function of the MPC; Specifically, $J_{1}$ is the PS field, which serves as a penalty as robot enters human’s personal space, where, $N_{human}$ denotes the number of humans in the environment and $PS(q)$ represents the social cost associated with the $q$th human; $J_{2}$ penalizes the tracking error between the predicted output $\eta$ and $\eta_{\text{Ref}}$ throughout the prediction horizon, $N_{p}$, weighted by the positive definite matrix $Q$; $J_{3}$ regulates the front and rear wheels’ steering action $\delta=\begin{bmatrix}\delta_{f}&\delta_{r}\end{bmatrix}^{T}$ and the deviation of the robot front and rear wheels’ velocity $V=\begin{bmatrix}v_{f}&v_{r}\end{bmatrix}^{T}$ from its reference $V_{ref}=\begin{bmatrix}v_{f,ref}&v_{r,ref}\end{bmatrix}^{T}$ over the control horizon $N_{c}$, with weighting matrices $R$ and $T$, respectively.
Furthermore, Equ. 21 and 22 describe the motion model of the robot by using the kinematic model as in the discrete time domain, where $A_{k}$, $B_{k}$, and $C_{k}$ are the linearized system matrices at time step $k$, and $d_{k}$ represents model mismatch terms compared with the nonlinear model. Constraints are imposed on both the control increment and the absolute control input using Equ. 23 and Equ. 24, where $\Delta u_{c\min}$ and $\Delta u_{c\max}$ denote the lower and upper bounds of control rate changes, and $u_{c\min}$ and $u_{c\max}$ define the control limits. In addition, the inequality Equ. 25 is a linearized wheel-speed constraint designed to prevent slippage due to speed differences between the front and rear wheels in the body-fixed frame, where $g(u_{c})=|v_{f}cos(\delta_{f})-v_{r}cos(\delta_{r})|$ is the nonlinear wheel difference expression; $E(u_{c})$ is the grandiant of the function $g$ with respect of $u_{c}$. Without this constraint, the motion planner may generate control inputs that violate the non-slip assumption in the kinematic motion model, potentially leading to unreliable or unsafe robot behaviour. Additional details regarding this constraint can be found in [40]. Finally, Equ. 26 specifies the initial condition of the optimization problem as the system states at the current time step.
$\arg\min_{u_{c}}J=J_{1}+J_{2}+J_{3}$ (18)
@@ V-C Model Predictive Control-based Planner Design
$J_{1}=\sum_{q=1}^{N_{human}}PS(q),J_{2}=\sum_{i=0}^{N_{p}-1}\|\eta(k+i+1)-\eta_{\text{Ref}}\|_{Q}^{2}$ (19)
$J_{3}=\sum_{j=0}^{N_{c}-1}\|\delta(k+j)\|^{2}_{R}+\|V(k+j)-V_{\text{Ref}}\|_{T}^{2}$ (20)
@@
$\zeta(k+i+1)=A_{k}\zeta(k+i)+B_{k}u_{c}(k+j)+d_{k}(k+i)$ (21)
$\eta(k+i+1)=C_{k}\zeta(k+i+1)$ (22)
@@
$\left|E(u_{c}(0))u_{c}(k+j)+g(u_{c}(0))\right|\leq 0.1$ (25)
@@
VI Experiments
In this section, we present experimental results to evaluate the effectiveness and practicality of the proposed SAP-CoPE framework. We first validate the perception system to verify that its detection accuracy compares with baseline methods, and also in the presence of computation and communication delays. We then evaluate the overall system across a range of scenarios with different numbers of humans, with focus on the motion planning system and comparison with other planning algorithms.
VI-A Perception Experiments
VI-A1 Evaluation Metrics
@@
Figure 5: Spatial distribution of position errors.
Figure 6: Spatial distribution of yaw errors.
TABLE II: Pose Estimation Performance Evaluation